# BiT-MamSleep: Bidirectional Temporal Mamba for EEG Sleep Staging

### v1

@@ IV-D Ablation Study
  • Bidirectional Mamba (BiMamba): This is the full MamSleep model, where both forward and backward temporal dependencies are modeled simultaneously, providing a comprehensive representation of EEG signals.
Each of these experiments was evaluated using common metrics including Accuracy (ACC), Macro-averaged F1-score (MF1), and Cohen’s Kappa ($\kappa$). The results of the ablation experiments are summarized in Table 4.
The ablation study results demonstrate that the Bidirectional Mamba (BiMamba) mechanism significantly enhances the MamSleep model’s performance, outperforming other configurations. Without BiMamba, the model’s accuracy, MF1, and Cohen’s Kappa values decreased, highlighting the limitations of capturing only spatial features. Using only forward Mamba showed some improvement over the baseline (TRCNN + AFMs), indicating the value of capturing past dependencies. However, it still lagged behind BiMamba, as forward modeling alone is insufficient for capturing complete temporal context needed for accurate sleep stage transitions. Similarly, using only backward Mamba improved performance compared to the baseline but was still inferior to the full BiMamba model, due to the lack of past context. The BiMamba’s bidirectional approach effectively captures both past and future dependencies, leading to substantial improvements, particularly in handling imbalanced classes and subtle sleep stage transitions, making it the most effective approach for accurate sleep stage classification.
@@ V Conclusion
In this paper, we presented MamSleep, a novel architecture designed for automatic sleep stage classification using EEG data. Our model integrates the Triple-Resolution CNN (TRCNN) for efficient multi-scale feature extraction and the Bidirectional Mamba (BiMamba) mechanism, which effectively captures both short- and long-term temporal dependencies through bidirectional processing. The experimental results on multiple public datasets demonstrate that MamSleep significantly outperforms state-of-the-art methods in terms of accuracy, F1-score, and Kappa, particularly excelling in handling long EEG sequence.
@@
While MamSleep demonstrates significant improvements, future work could explore optimizing the computational complexity of the model, making it more suitable for real-time applications such as wearable devices for sleep monitoring. Additionally, we plan to evaluate the generalizability of MamSleep on other physiological signal datasets and investigate its potential for broader health monitoring applications. By advancing EEG-based sleep classification, MamSleep aims to contribute to more accessible and accurate sleep health assessments in clinical and consumer settings.

### v2

@@ IV-D Ablation Study
  • BiMamba: This is the full MamSleep model, where both forward and backward temporal dependencies are modeled simultaneously, providing a comprehensive representation of EEG signals.
Each of these experiments was evaluated using common metrics including ACC, MF1, and $\kappa$. The results of the ablation experiments are summarized in Table 4.
The ablation study results demonstrate that the BiMamba mechanism significantly enhances the BiT-MamSleep’s performance, outperforming other configurations. Without BiMamba, the model’s accuracy, MF1, and Cohen’s Kappa values decreased, highlighting the limitations of capturing only spatial features. Using only forward Mamba showed some improvement over the baseline (TRCNN + AFMs), indicating the value of capturing past dependencies. However, it still lagged behind BiMamba, as forward modeling alone is insufficient for capturing complete temporal context needed for accurate sleep stage transitions. Similarly, using only backward Mamba improved performance compared to the baseline but was still inferior to the full BiMamba model, due to the lack of past context. The BiMamba’s bidirectional approach effectively captures both past and future dependencies, leading to substantial improvements, particularly in handling imbalanced classes and subtle sleep stage transitions, making it the most effective approach for accurate sleep stage classification.
@@ V Conclusion
In this paper, we presented MamSleep, a novel architecture designed for automatic sleep stage classification using EEG data. Our model integrates the TRCNN for efficient multi-scale feature extraction and the BiMamba mechanism, which effectively captures both short- and long-term temporal dependencies through bidirectional processing. The experimental results on multiple public datasets demonstrate that MamSleep significantly outperforms state-of-the-art methods in terms of accuracy, F1-score, and Kappa, particularly excelling in handling long EEG sequence.
@@
While BiT-MamSleep demonstrates significant improvements, future work could explore optimizing the computational complexity of the model, making it more suitable for real-time applications such as wearable devices for sleep monitoring. Additionally, we plan to evaluate the generalizability of BiT-MamSleep on other physiological signal datasets and investigate its potential for broader health monitoring applications. By advancing EEG-based sleep classification, BiT-MamSleep aims to contribute to more accessible and accurate sleep health assessments in clinical and consumer settings.